# The Low-Degree Hardness of Finding Large Independent Sets in Sparse Random Hypergraphs

### v1

@@ Abstract
We study the algorithmic task of finding large independent sets in sparse Erdős–Rényi random $r$-uniform hypergraphs on $n$ vertices having average degree $d$. Krivelevich and Sudakov showed that the maximum independent set has density $\left(\frac{r}{r-1}\cdot\frac{\log d}{d}\right)^{1/(r-1)}$ in the double limit $n\to\infty$ followed by $d\to\infty$. We show that the class of low-degree polynomial algorithms can find independent sets of density $\left(\frac{1}{r-1}\cdot\frac{\log d}{d}\right)^{1/(r-1)}$ but no larger. This extends and generalizes earlier results of Gamarnik and Sudan, Rahman and Virág, and Wein on graphs, and answers a question of Bal and Bennett. We conjecture that this statistical-computational gap of a multiplicative factor of $r^{-1/(r-1)}$ indeed holds for this problem.
@@
While the graph case has been extensively studied, this work is the first to consider statistical-computational gaps of optimization problems on random hypergraphs. Our results suggest that these gaps persist for larger uniformities as well as across many models. A somewhat surprising aspect of the gap for balanced independent sets is that the algorithm achieving the lower bound is a simple degree-$1$ polynomial.
@@ 1 Introduction
The problem of finding a maximum size independent set in hypergraphs is of fundamental interest, both in practical and theoretical aspects. It arises in various applications in data mining, image processing, database design, and parallel computing, to name a few. It is well-known that finding the maximum independent set in graphs is an NP-hard problem [Kar10]. In fact, it is hard to approximate the maximum independent set in a graph to within $n^{1-\varepsilon}$ for any $\varepsilon>0$ [Has96, Kho01, Zuc06]. As hypergraphs are a generalization of graphs, these hardness results clearly extend to this setting. In order to better understand the hardness of maximum independent set, a natural line of research is to analyze the structure of the “hard” instances. We consider a somewhat related question: is maximum independent set tractable on “typical” instances? In particular, we consider the $r$-uniform Erdős–Rényi hypergraphs $\mathcal{H}_{r}(n,p)$, where each potential edge in an $n$-vertex $r$-uniform hypergraph is included independently with probability $p$ (see Definition 1.4 for a formal description).
@@
The study of finding large independent sets in random graphs dates back to Matula [Mat76] who showed that an Erdős–Rényi graph does not contain an independent set of size at least $(2+o(1))\log_{1/(1-p)}n$ with high probability for a large range of $p$. A simple algorithm can construct an independent set of size $\sim\log_{1/(1-p)}n$, however, finding a larger one seems to be hard. In 1976, Karp conjectured that no polynomial time algorithm can compute an independent set of size at least $(1+\varepsilon)\log_{2}n$ for any $\varepsilon>0$ with high probability for $p=1/2$ [Kar76]. This conjecture has spurred a large research effort in order to determine the so-called statistical-computational gap for a number of “hard” problems on random graphs (see for example [Jer92, DM15, GJW20, Wei22, BH22]).
@@
We are interested in the sparse regime, i.e., $p=d/\binom{n-1}{r-1}$, where $d$ is a constant. Note that for our choice of $p$, the expected degree of a vertex is precisely $d$. For $r=2$, Frieze showed that with high probability as $n\to\infty$ the maximum independent set of $\mathcal{H}_{2}(n,p)$ has density $(2\pm o_{d}(1))\log d/d$ for sufficiently large constant $d$ [Fri90]. It is conjectured that no polynomial time algorithm can find an independent set of density $(1+\varepsilon)\log d/d$ with high probability for any $\varepsilon>0$, i.e., there is a statistical-computational gap of a multiplicative factor of $1/2$. Proving this conjecture is equivalent to proving a statement stronger than $P=NP$ and so researchers have focused on providing evidence of intractability through a variety of methods. One such method is to prove intractability for restricted classes of algorithms (see §1.1 for examples). Gamarnik and Sudan [GS14] first proved a gap of a multiplicative factor of $(1/2+1/2\sqrt{2})$ for local algorithms and pioneered the framework of the Overlap Gap Property, which subsequently led to proving the conjectured gap of a multiplicative factor of $1/2$ for both local algorithms [RV17] and low-degree algorithms [Wei22]. In our work we focus on low-degree algorithms (see §1.2 for an overview of the framework).
@@
Moving beyond graphs to hypergraphs, Krivelevich and Sudakov showed that with high probability the maximum independent set of $\mathcal{H}_{r}(n,p)$ has density $(1\pm o(1))\left(\frac{r}{r-1}\cdot\frac{\log d}{d}\right)^{1/(r-1)}$ [KS98], which matches Frieze’s result when $r=2$. It is natural to ask whether this phenomenon of a statistical-computational gap for finding large independent sets extends to the hypergraph setting. Additionally, determining the size of this gap is of interest as well. To the best of our knowledge, this paper is the first to consider the computational threshold for independent sets in random hypergraphs.
@@
As local algorithms are a subclass of low-degree algorithms, our results answer a question of Bal and Bennett [BB23, §6.2]. In light of this result and the aforementioned statistical-computational gap conjecture of independent sets in random graphs, we make the following analogous conjecture for hypergraphs.
@@
In recent work, Perkins and the second author explored the universality of the statistical-computati-
@@
onal gap of independent sets in graphs [PW24]. They considered the following question: in what other random graph models (and for what kinds of independent sets) do we see such a gap? Specifically, they studied independent sets of the Erdős–Rényi bipartite graph with $n$ vertices in each partition. When considering bipartite graphs, there is a trivial independent set of density at least $1/2$, namely, the larger partition. Furthermore, there exists a max-flow based polynomial time algorithm to find the maximum independent set in bipartite graphs. As observed by Perkins and the second author, imposing global constraints on the independent set $I$ can introduce computational intractability.
@@
When considering bipartite graphs, a natural global constraint to consider is balancedness. Let $G=(V_{1}\cup V_{2},\,E)$ be a bipartite graph. We say an independent set $I\subseteq V(G)$ is balanced if $|I\cap V_{1}|=|I|/2$, i.e., $I$ contains an equal number of vertices from each partition. There has been extensive research in determining the density of the maximum balanced independent set in both deterministic and random bipartite graphs [FMO93, Axe+21, Cha23]. Perkins and the second author showed that the Erdős–Rényi bipartite graph with $n$ vertices in each partition has a balanced independent set with density $(2\pm o(1))\log d/d$ with high probability. They also showed that the low-degree computational threshold of the density of the maximum balanced independent set is $\log d/d$. In particular, balanced independent sets in bipartite graphs exhibit the same behavior as independent sets in graphs. We remark that Perkins and the second author prove a more general statement for so-called $\gamma$-balanced independent sets (see Definition 1.7).
@@
We extend these results to $r$-uniform $r$-partite hypergraphs. We say $H=(V,E)$ is $r$-partite if there exists a partition $V=V_{1}\cup\cdots\cup V_{r}$ such that each edge $e\in E$ contains precisely one vertex from each set $V_{i}$. Furthermore, if $|V_{1}|=\cdots=|V_{r}|$, we call such hypergraphs balanced. Multipartite hypergraphs find a wide array of applications in satisfiability problems, Steiner triple systems, and particle tracking in physics, to name a few. Note that the union of any $r-1$ partitions forms an independent set in such a hypergraph, and so the problem of finding an independent set of density at least $1-1/r$ is trivial. As this class of hypergaphs is a natural generalization of bipartite graphs, we may extend the definition of balanced independent sets to this setting. In particular, we say an independent set $I$ in an $r$-uniform $r$-partite hypergraph $H=(V_{1}\cup\cdots\cup V_{r},\,E)$ is balanced if $|I\cap V_{i}|=|I|/r$ for each $1\leq i\leq r$ ($I$ contains an equal number of vertices from each partition). Such a notion was introduced in recent work of the first author [Dha23], where he considered the size of the largest balanced independent set in deterministic $r$-partite hypergraphs.
@@
We consider the Erdős–Rényi $r$-uniform $r$-partite balanced hypergraph $\mathcal{H}(r,n,p)$ with $n$ vertices in each partition (note the notational distinction between $\mathcal{H}(r,n,p)$ and $\mathcal{H}_{r}(n,p)$). The second main result of this paper concerns the statistical-computational gap of finding large balanced independent sets in $\mathcal{H}(r,n,p)$ for $p=d/n^{r-1}$. We first prove a high probability bound on the asymptotic density of the largest balanced independent set in $\mathcal{H}(r,n,p)$. Moreover, we describe a simple, efficient algorithm that can find a balanced independent set within a multiplicative factor of $r^{-1/(r-1)}$ of the statistical threshold. This begs the following question: can one do better or is there a universality of the statistical–computational gap for finding large independent sets in hypergraphs as well? We provide evidence toward the latter.
@@
Our results answer a question of the first author, and for $r=2$, we recover the results of Perkins and the second author. We also conjecture that this statistical-computational gap persists for polynomial-time algorithms (a version of this conjecture for $r=2$ appeared in [PW24]).
@@
We do in fact prove more general statements for $\bm{\gamma}$-balanced independent sets (see Definition 1.7 for a formal definition). Here, rather than considering an independent set $I$ with an equal proportion of vertices in each partition, we specify the proportion of vertices in $I$ within each partition in a vector $\bm{\gamma}\in(0,1)^{r}$. For $\bm{\gamma}=(1/r,\ldots,1/r)$, we recover the original notion of balanced independent sets.
@@ 1.1 Relation to Prior Work
Inference problems with conjectured statistical-computational gaps are ubiquitous throughout statistics and computer science. Some classical examples include the planted clique problem [Jer92, DM15, MPW15, Bar+19] and community detection in random graphs [Dec+11, AV14, HS17], structured principal component analysis of matrices [BR13, LKZ15] and tensors [HSS15, Hop+17], and solving or refuting constraint satisfaction problems [AC08, Kot+17].
While proving statistical-computational gaps of the above problems is considered a daunting task, it has spurred a large research effort toward providing rigorous evidence of hardness for average-case instances. As mentioned earlier, this involves proving hardness for certain restricted classes of algorithms. These techniques include Markov chain Monte Carlo methods [Jer92, DFJ02], local algorithms [GS14, RV17, Che+19], belief propagation and approximate message passing algorithms [Dec+11, LKZ15], reductions from the presumed hard planted clique problem [HWX15, BB19], and statistical query models [Kea98, FPV15], to name a few.
Within the random graph framework, there are three primary problems of interest: optimization, testing, and estimation. This work focuses on the optimization problem, i.e., finding a large (balanced) independent set in a random hypergraph. One can analogously consider the testing and estimation problems by analyzing so-called “planted” models. The above techniques have been applied to a wide variety of such problems including the densest $k$-subgraph problem [FS97, Bha+10], planted dense subgraph problem [Jon+23, MWZ23, MWZ24], and hypergraph versions [Chl+18, Cor+22, DMW23].
@@
The main focus of this work is to provide sharp thresholds for the tractability of low-degree polynomial algorithms (to be formally defined in the next section). Here, each vertex’s membership in the (balanced) independent set is determined by a multivariate polynomial in the edge indicator variables of the hypergraph. This is a powerful class of algorithms as it includes the class of local algorithms as well as the algorithmic paradigms of approximate message passing and power iteration (see the discussion in [GJW24, Appendix A]). Furthermore, it has been observed that low-degree algorithms are as powerful as the best known polynomial-time algorithms for a number of problems in high-dimensional statistics including those mentioned in the preceding paragraphs.
There is by now a standard method for proving low-degree optimization bounds based on the Overlap Gap Property, which we describe in §1.4 (see the survey [Gam21] for a more extensive overview). This technique has been employed in a number of inference problems including independent sets in random graphs [GJW20, Wei22], random $k$-SAT [BH22], discrepancy of random matrices [Ven22], and random CSPs [CHM23]. Our work is the first to apply this strategy to random hypergraphs.
@@
Random graph inference has been a central research area for decades, while the hypergraph extension is less well-studied. In general, the hypergraph setting is believed to be considerably harder. In recent years, however, there has been an increase in both theoretical and applied interest in this setting. These include applying spectral methods to solve testing and estimation problems [LZ22, Jon+23] and determining statistical and computational thresholds for planted problems [YS21, YS21a, DMW23].
While the maximum independent set problem is NP-hard, there has been progress for restricted classes of algorithms on structurally constrained hypergraphs. In particular, [HL09, GS11] consider bounded-degree hypergraphs, [Hal+16] consider streaming algorithms for sparse hypergraphs, [KLP21] consider semi-random hypergraphs, and [Hal02, AHL13] explore SDP based methods of solving maximum independent set. To the best of our knowledge, this is the first paper to consider the statistical-computational gap of finding independent sets in Erdős–Rényi hypergraphs.
In contrast, multipartite hypergraphs have been less extensively studied from a computational standpoint, however, there are a number of theoretical results which may be of interest to the reader [KSV17, Dha23a, BM24]. In [GSS15], the authors consider the minimum vertex cover problem on such hypergraphs. A somewhat surprising result is that while the problem is tractable for $r=2$, it is NP-hard for $r\geq 3$. This makes our results all the more interesting as we show that the behavior of maximum balanced independent set does not exhibit such a distinction based on $r$. A similar observation was made in [BWZ12] who consider the appearance of a $k$-core (a subhypergraph of minimum degree $k$) in certain random $r$-partite hypergraph models.

### v2

@@ Abstract
We study the algorithmic task of finding large independent sets in sparse Erdős–Rényi random $r$-uniform hypergraphs on $n$ vertices having average degree $d$. Krivelevich and Sudakov showed that the maximum independent set has density $\left(\frac{r}{r-1}\cdot\frac{\log d}{d}\right)^{1/(r-1)}$ in the double limit $n\to\infty$ followed by $d\to\infty$. We show that the class of low-degree polynomial algorithms can find independent sets of density $\left(\frac{1}{r-1}\cdot\frac{\log d}{d}\right)^{1/(r-1)}$ but no larger. This extends and generalizes earlier results of Gamarnik and Sudan, Rahman and Virág, and Wein on graphs, and answers a question of Bal and Bennett. We conjecture that this statistical–computational gap of a multiplicative factor of $r^{-1/(r-1)}$ indeed holds for this problem.
@@
While the graph case has been extensively studied, this work is the first to consider statistical–computational gaps of optimization problems on random hypergraphs. Our results not only pin down the precise threshold for low-degree algorithms in two different settings, but also suggest that these gaps persist for larger uniformities as well as across many models. A somewhat surprising aspect of the gap for balanced independent sets is that the algorithm achieving the lower bound is a simple degree-$1$ polynomial.
@@ 1 Introduction
The problem of finding a maximum size independent set in hypergraphs is of fundamental interest, both in practical and theoretical aspects. It arises in various applications in data mining, image processing, database design, and parallel computing, to name a few. It is well-known that finding the maximum independent set in graphs is an NP-hard problem [karp2010reducibility]. In fact, it is hard to approximate the maximum independent set in a graph to within $n^{1-\varepsilon}$ for any $\varepsilon>0$ [hastad1996clique, khot2001improved, zuckerman2006linear]. As hypergraphs are a generalization of graphs, these hardness results clearly extend to this setting. In order to better understand the hardness of maximum independent set, a natural line of research is to analyze the structure of the “hard” instances. We consider a somewhat related question: is maximum independent set tractable on “typical” instances? In particular, we consider the $r$-uniform Erdős–Rényi hypergraphs $\mathcal{H}_{r}(n,p)$, where each potential edge in an $n$-vertex $r$-uniform hypergraph is included independently with probability $p$ (see Definition 1.4 for a formal description).
@@
The study of finding large independent sets in random graphs dates back to Matula [matula1976largest] who showed that an Erdős–Rényi graph does not contain an independent set of size at least $(2+o(1))\log_{1/(1-p)}n$ with high probability for a large range of $p$. A simple algorithm can construct an independent set of size $\sim\log_{1/(1-p)}n$, however, finding a larger one seems to be hard. In 1976, Karp conjectured that no polynomial time algorithm can compute an independent set of size at least $(1+\varepsilon)\log_{2}n$ for any $\varepsilon>0$ with high probability for $p=1/2$ [karp1976probabilistic]. This conjecture has spurred a large research effort in order to determine the so-called statistical–computational gap for a number of “hard” problems on random graphs (see for example [jerrum1992large, deshpande2015improved, gamarnik2020low, wein2022optimal, bresler2022algorithmic]).
@@
We are interested in the sparse regime, i.e., $p=d/\binom{n-1}{r-1}$, where $d$ is a constant. Note that for our choice of $p$, the expected degree of a vertex is precisely $d$. For $r=2$, Frieze showed that with high probability as $n\to\infty$ the maximum independent set of $\mathcal{H}_{2}(n,p)$ has density $(2\pm o_{d}(1))\log d/d$ for sufficiently large constant $d$ [frieze1990independence]. It is conjectured that no polynomial time algorithm can find an independent set of density $(1+\varepsilon)\log d/d$ with high probability for any $\varepsilon>0$, i.e., there is a statistical–computational gap of a multiplicative factor of $1/2$. Proving this conjecture is equivalent to proving a statement stronger than $P=NP$ and so researchers have focused on providing evidence of intractability through a variety of methods. One such method is to prove intractability for restricted classes of algorithms (see §1.1 for examples). Gamarnik and Sudan [gamarnik2014limits] first proved a gap of a multiplicative factor of $(1/2+1/2\sqrt{2})$ for local algorithms and pioneered the framework of the Overlap Gap Property, which subsequently led to proving the conjectured gap of a multiplicative factor of $1/2$ for both local algorithms [rahman2017local] and low-degree algorithms [wein2022optimal]. In our work we focus on low-degree algorithms (see §1.2 for an overview of the framework).
@@
Moving beyond graphs to hypergraphs, Krivelevich and Sudakov showed that with high probability the maximum independent set of $\mathcal{H}_{r}(n,p)$ has density $(1\pm o(1))\left(\frac{r}{r-1}\cdot\frac{\log d}{d}\right)^{1/(r-1)}$ [krivelevich1998chromatic], which matches Frieze’s result when $r=2$. It is natural to ask whether this phenomenon of a statistical–computational gap for finding large independent sets extends to the hypergraph setting. Additionally, determining the size of this gap is of interest as well. To the best of our knowledge, this paper is the first to consider the computational threshold for independent sets in random hypergraphs.
@@
As local algorithms are a subclass of low-degree algorithms, our results answer a question of Bal and Bennett [bal2023larger, §6.2]. In light of this result and the aforementioned statistical–computational gap conjecture of independent sets in random graphs, we make the following analogous conjecture for hypergraphs.
@@
In recent work, Perkins and the second author explored the universality of the statistical–computational gap of independent sets in graphs [perkins2024hardness]. They considered the following question: in what other random graph models (and for what kinds of independent sets) do we see such a gap? Specifically, they studied independent sets of the Erdős–Rényi bipartite graph with $n$ vertices in each partition. When considering bipartite graphs, there is a trivial independent set of density at least $1/2$, namely, the larger partition. Furthermore, there exists a max-flow based polynomial time algorithm to find the maximum independent set in bipartite graphs. As observed by Perkins and the second author, imposing global constraints on the independent set $I$ can introduce computational intractability.
@@
When considering bipartite graphs, a natural global constraint to consider is balancedness. Let $G=(V_{1}\cup V_{2},\,E)$ be a bipartite graph. We say an independent set $I\subseteq V(G)$ is balanced if $|I\cap V_{1}|=|I|/2$, i.e., $I$ contains an equal number of vertices from each partition. There has been extensive research in determining the density of the maximum balanced independent set in both deterministic and random bipartite graphs [favaron1993bipartite, axenovich2021bipartite, chakraborti2023extremal]. Perkins and the second author showed that the Erdős–Rényi bipartite graph with $n$ vertices in each partition has a balanced independent set with density $(2\pm o(1))\log d/d$ with high probability. They also showed that the low-degree computational threshold of the density of the maximum balanced independent set is $\log d/d$. In particular, balanced independent sets in bipartite graphs exhibit the same behavior as independent sets in graphs. We remark that Perkins and the second author prove a more general statement for so-called $\bm{\gamma}$-balanced independent sets (see Definition 1.7).
@@
We extend these results to $r$-uniform $r$-partite hypergraphs. We say $H=(V,E)$ is $r$-partite if there exists a partition $V=V_{1}\cup\cdots\cup V_{r}$ such that each edge $e\in E$ contains precisely one vertex from each set $V_{i}$. Furthermore, if $|V_{1}|=\cdots=|V_{r}|$, we call such hypergraphs balanced. Multipartite hypergraphs find a wide array of applications in satisfiability problems, Steiner triple systems, and particle tracking in physics, to name a few. Note that the union of any $r-1$ partitions forms an independent set in such a hypergraph, and so the problem of finding an independent set of density at least $1-1/r$ is trivial. As this class of hypergaphs is a natural generalization of bipartite graphs, we may extend the definition of balanced independent sets to this setting. In particular, we say an independent set $I$ in an $r$-uniform $r$-partite hypergraph $H=(V_{1}\cup\cdots\cup V_{r},\,E)$ is balanced if $|I\cap V_{i}|=|I|/r$ for each $1\leq i\leq r$, i.e., $I$ contains an equal number of vertices from each partition. Such a notion was introduced in recent work of the first author [dhawan2023balanced], who considered the size of the largest balanced independent set in deterministic $r$-partite hypergraphs.
@@
We consider the Erdős–Rényi $r$-uniform $r$-partite balanced hypergraph $\mathcal{H}(r,n,p)$ with $n$ vertices in each partition (note the notational distinction between $\mathcal{H}(r,n,p)$ and $\mathcal{H}_{r}(n,p)$). The second main result of this paper concerns the statistical–computational gap of finding large balanced independent sets in $\mathcal{H}(r,n,p)$ for $p=d/n^{r-1}$. We first prove a high probability bound on the asymptotic density of the largest balanced independent set in $\mathcal{H}(r,n,p)$. Moreover, we describe a simple, efficient algorithm that can find a balanced independent set within a multiplicative factor of $r^{-1/(r-1)}$ of the statistical threshold. This begs the following question: can one do better or is there a universality of the statistical–computational gap for finding large independent sets in hypergraphs as well? We provide evidence toward the latter.
@@
Our results answer a question of the first author, and for $r=2$, we recover the results of Perkins and the second author. We also conjecture that this statistical–computational gap persists for polynomial-time algorithms (a version of this conjecture for $r=2$ appeared in [perkins2024hardness]).
@@
We do in fact prove more general statements for $\bm{\gamma}$-balanced independent sets. Here, rather than considering an independent set $I$ with an equal proportion of vertices in each partition, we specify the proportion of vertices in $I$ within each partition in a vector $\bm{\gamma}\in(0,1)^{r}$ (see Definition 1.7 for a formal definition). For $\bm{\gamma}=(1/r,\ldots,1/r)$, we recover the original notion of balanced independent sets.
@@ 1.1 Relation to Prior Work
Inference problems with conjectured statistical–computational gaps are ubiquitous throughout statistics and computer science. Some classical examples include the planted clique problem [jerrum1992large, deshpande2015improved, meka2015sum, barak2019nearly] and community detection in random graphs [decelle2011asymptotic, arias2014community, hopkins2017bayesian], structured principal component analysis of matrices [berthet2013computational, lesieur2015phase] and tensors [hopkins2015tensor, hopkins2017power], and solving or refuting constraint satisfaction problems [achlioptas2008algorithmic, kothari2017sum].
While proving statistical–computational gaps of the above problems is considered a daunting task, it has spurred a large research effort toward providing rigorous evidence of hardness for average-case instances. As mentioned earlier, this involves proving hardness for certain restricted classes of algorithms. These techniques include Markov chain Monte Carlo methods [jerrum1992large, dyer2002counting], local algorithms [gamarnik2014limits, rahman2017local, chen2019suboptimality], belief propagation and approximate message passing algorithms [decelle2011asymptotic, lesieur2015phase], reductions from the presumed hard planted clique problem [hajek2015computational, brennan2019optimal], and statistical query models [kearns1998efficient, feldman2015complexity], to name a few.
Within the random graph framework, there are three primary problems of interest: optimization, testing, and estimation. This work focuses on the optimization problem, i.e., finding a large (balanced) independent set in a random hypergraph. One can analogously consider the testing and estimation problems by analyzing so-called “planted” models. The above techniques have been applied to a wide variety of such problems including the densest $k$-subgraph problem [feige1997densest, bhaskara2010detecting], planted dense subgraph problem [jones2023sum, PDC, PDC_IT], and hypergraph versions [chlamtac2018densest, corinzia2022statistical, dhawan2023detection].
@@
The study of low-degree algorithms has become a popular approach in recent years. The main focus of this work is to provide sharp thresholds for the tractability of low-degree polynomial algorithms (to be formally defined in the next section). Here, each vertex’s membership in the (balanced) independent set is determined by a multivariate polynomial in the edge indicator variables of the hypergraph. This is a powerful class of algorithms as it includes the class of local algorithms as well as the algorithmic paradigms of approximate message passing and power iteration (see the discussion in [gamarnik2024hardness, Appendix A]). Given the current state of average-case complexity theory, it has been shown that low-degree algorithms are as powerful as the best known polynomial-time algorithms for a number of problems in high-dimensional statistics including those mentioned in the preceding paragraphs. Therefore, the failure of low-degree algorithms is a sign of concrete evidence for computational hardness of statistical problems.
There is by now a standard method for proving low-degree optimization bounds based on the Overlap Gap Property, which we describe in §1.4 (see the survey [gamarnik2021overlap] for a more extensive overview). This technique has been employed in a number of inference problems including independent sets in random graphs [gamarnik2020low, wein2022optimal], random $k$-SAT [bresler2022algorithmic], discrepancy of random matrices [venkat2022efficient], and random CSPs [chen2023local]. Our work is the first to apply this strategy to random hypergraphs.
@@
Random graph inference has been a central research area for decades, while the hypergraph extension is less well-studied. In general, the hypergraph setting is believed to be considerably harder. In recent years, however, there has been an increase in both theoretical and applied interest in this setting. These include applying spectral methods to solve testing and estimation problems [luo2022tensor, jones2023sum] and determining statistical and computational thresholds for planted problems [yuan2021heterogeneous, yuan2021information, dhawan2023detection].
While the maximum independent set problem is NP-hard, there has been progress for restricted classes of algorithms on structurally constrained hypergraphs. In particular, [halldorsson2009independent, guruswami2011complexity] consider bounded-degree hypergraphs, [halldorsson2016streaming] consider streaming algorithms for sparse hypergraphs, [khanna2021independent] consider semi-random hypergraphs, and [halperin2002improved, agnarsson2013sdp] explore SDP based methods of solving maximum independent set. To the best of our knowledge, this is the first paper to consider the statistical–computational gap of finding independent sets in Erdős–Rényi hypergraphs.
In contrast, multipartite hypergraphs have been less extensively studied from a computational standpoint, however, there are a number of theoretical results which may be of interest to the reader [kamvcev2017bounded, dhawan2023list, bowtell2024matchings]. In [guruswami2015inapproximability], the authors consider the minimum vertex cover problem on such hypergraphs. A somewhat surprising result is that while the problem is tractable for $r=2$, it is NP-hard for $r\geq 3$. This makes our results all the more interesting as we show that the behavior of maximum balanced independent set does not exhibit such a distinction based on $r$. A similar observation was made in [botelho2012cores] who consider the appearance of a $k$-core (a subhypergraph of minimum degree $k$) in certain random $r$-partite hypergraph models.